### v1

# Inflection Point of Minimally Coupled Tachyonic Scalar Field

Abstract
In this paper, we investigate the behaviour of a minimally coupled tachyonic scalar field at inflection points in an accelerating universe. We consider the different expansion factors and obtain potentials of tachyonic scalar field. Inflection points of homogeneous tachyonic scalar field is calculated for these potentials. We employ the tools of dynamical system analysis for the considered potentials and obtain the stable points.
1 Introduction
The cosmic acceleration is now an established phenomenon. As supported by several cosmological and astrophysical observations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] our universe is in the phase of accelerated expansion at the present stage and generally, it is believed that behind this accelerated expansion there is a mysterious component with negative pressure dominating $70\%$ of the universe. One of the good candidates that can explain this acceleration is ”Dark Energy” [11, 12, 13, 14, 15] which is introduced as modified matter in Einstein’s field equation which sits with the energy-momentum tensor. One of the remarkable features of dark energy is its negative pressure and repulsive gravity characteristics led by the equation of state $w_{de}<-1/3$. In alternative gravity theories, cosmic acceleration can be accounted for by modifications to the geometrical aspects of the field equation. Observationally, the data from Planck-2018 CMB [7, 8, 9, 10] proposes that $70\%$ of the energy budget in the universe is in the form of dark energy while the remaining $30\%$ of the energy budget of the universe is attributed to non-relativistic baryonic and non-baryonic pressureless ( $p_{m}=0$) dust matter with equation of state(EoS) $w_{m}=0$. In recent years, various candidates have been put forth to elucidate the nature of dark energy, with the cosmological constant standing out as one of the simplest models, embodying constant dark energy with an EoS of $w_{\Lambda}=-1$. Despite its alignment with observational data to a certain accuracy, the $\Lambda$CDM model grapples with unresolved cosmological constant and coincidence problems [16, 17, 18, 19, 20]. In response to these limitations, dynamical dark energy models, particularly those involving scalar fields, have gained prominence [21, 22, 23, 24, 25, 26] and with scalar field serving as a candidate of dynamical dark energy dubbed as $\phi$CDM having dynamical EoS $w_{\phi}$.
The nature of dark energy has been the subject of several literature; as alternatives to the cosmological constant, dynamic dark energy models have been developed, in which the EoS varies with time. There are many different dynamical dark energy models like quintessence field having tracker behaviour, late time accelerating solutions of modified Friedmann equations lead by modification of gravity, some other models include tachyon models, holographic dark energy models, chaplygin gas, phantom and k-essence [28, 29, 30, 31, 32, 33, 34, 35].
In our paper, we have focused on the tachyonic field as a candidate for dark energy. A scalar field of this kind is driven by string theory as the negative-mass mode of the open string perturbative spectrum, though the primary use of it in the dark energy sector is phenomenological here in our paper, we have investigated the behaviour if tachyonic scalar field near infection point. Recently, the dark energy models with the inflection point in the context of inflation have been studied by several authors [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
To obtain a contemporary value of EoS $w_{\phi}\approx-1$ i.e. to get accelerated expansion mechanism of slow rolling is the key ingredient so, by imposing the approximation in which the kinetic term $\dot{\phi}^{2}/2$ of tachyon is significantly smaller than the potential $V(\phi)$ at the present moment i.e. $\dot{\phi}^{2}/2<<V(\phi)$ and this can be obtained by sitting $\phi$ in an extremely flat region of the potential.We consider the late time behavior of Tachyon scalar field in the accelerated expansion of the universe, where few similarities with early inflation. Numerous authors have examined the cosmic behaviour of the tachyonic scalar field [57, 58, 59], discovering possibilities for both pure exponential growth and power law expansion. We try to investigate the possibilities based on the superposition of these forms because the current data have not yet been able to determine the precise shape of the evolution of the scale factor.
1.1 Dynamical set up
In this section, we present the mathematical framework on which the dynamics are based. We choose GR to be the background theory. The cosmological principle is encoded in the FLRW metric when expressed in terms of spherical co-moving coordinates looks
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}\equiv-dt^{2}+a^{2}(t)\left(\frac{dr^{2}}{1-kr^{2}}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}\right).$ (1)
Here we have taken $c=1$. In the above equation, $a(t)$ is the scale factor and the parameter $k$ is the spatial curvature having values $k=-1,0,1$ corresponding to spatially open, flat, and closed geometry respectively. For our model, we have taken the effective four-dimensional action for the tachyonic field which is given as [62, 63]:
$S==\frac{1}{2\kappa^{2}}\int d^{4}xf(\phi)\sqrt{-g}R+\int d^{4}xV(\phi)\sqrt{-g}\sqrt{\left(1+g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right)},$ (2)
where $V(\phi)$ is the potential of tachyonic scalar field and $\kappa^{2}=\frac{1}{M^{2}_{p}}=1$ set the 4D Planck scale considering $M_{p}$ is Planck mass as well as function of tachyon field $f(\phi)=1$ under minimal coupling. One can obtain the energy-momentum tensor for this action as
$T^{(\phi)}_{\mu\nu}=\frac{V(\phi)\partial_{\mu}\phi\partial_{\nu}\phi}{\sqrt{1+g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi}}-g_{\mu\nu}V(\phi)\sqrt{1+g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi}$ (3)
Einstein’s field equation for a spatially flat universe ($k=0$) gives the Friedmann equations as:
$H^{2}=\left(\frac{\dot{a}}{a}\right)^{2}=\frac{\kappa^{2}}{3}\rho,$ (4)
and
$\frac{\ddot{a}}{a}=-\frac{\kappa^{2}}{2}(\rho+p),$ (5)
where $H(t)\equiv\frac{da}{dt}$ is the Hubble parameter, $\rho$ and $p$ are the energy and the pressure respectively. In general, there could be multiple components of matter contributing to energy-momentum tensor, each of which with a different equation of state. However, in the post-recombination era (after $z\longrightarrow 1000)$ and more effectively towards the present epoch $(z\longrightarrow 0)$ matter and dark energy dominate the universe. We are therefore driven to focus on only two components: matter and dark energy. These two components fit together so perfectly that, if one is matter $w_{m}\approx 0$, the other is the dark energy $(w_{de}\approx-1$). For cosmic acceleration, we have $\rho+3p<0$. The continuity equation is given as:
$\dot{\rho}+3H(1+w)\rho=0.$ (6)
In Sec.2 investigation of the behaviour of tachyonic scalar field is carried out near the inflection point. We have considered three different possibilities of cosmic behaviours; first two cases of conventional power-law expansion and pure exponential growth i.e. $a(t)=\alpha t^{n},a(t)=\gamma e^{\beta t}$ respectively and the third case is of quasi-exponential expansion of the scale factor $a(t)=\eta t^{n}e^{\beta t}$ for these three cases we have calculated the respective potentials and then we investigated the behaviour of that particular model near their respective inflection points. In Sec.3 dynamical analysis for the above two cases is carried out.
2 Inflection point of tachyonic scalar field:
Using Eq.(3) for flat ($K=0$) FRLW universe the energy density and pressure for the tachyonic field are given by:
$\rho_{\phi}=\frac{V(\phi)}{\sqrt{1-\dot{\phi}^{2}}},$ (7)
$p_{\phi}=-V(\phi)\sqrt{1-\dot{\phi}^{2}}.$ (8)
Here dot signifies the derivative with respect to time. $V(\phi)$ is the potential associated with the tachyonic field that we have calculated in the upcoming subsections for each case. Now by using Eq.(7) and Eq.(8) the model may be written as fluid. By using:
$\frac{\ddot{a}}{a}=H^{2}+\dot{H}^{2}$ (9)
and plugging Eq.(7) and Eq.(8) in Eq.(4) and (5) we get the following form of the second Friedmann equation:
$\frac{\ddot{a}}{a}=\frac{\kappa^{2}V(\phi)}{3\sqrt{1-\dot{\phi}^{2}}}\left(1-\frac{3}{2}\dot{\phi}^{2}\right)$ (10)
Now as we know for the accelerated expansion of the universe we have $\ddot{a}>0$ thus using Eq.10; $\dot{\phi}^{2}<\frac{2}{3}$ and the EoS (Equation of state) for $\phi$ is given by:
$w_{\phi}=\frac{p_{\phi}}{\rho_{\phi}}=(\dot{\phi}^{2}-1)$ (11)
For cosmic acceleration the possible range of the values of $w_{\phi}$ is $-1\leq w_{\phi}<-\frac{1}{3}$ and $0<\dot{\phi}^{2}<\frac{2}{3}$.
Using Eq.(7) and (8) one can easily find the equation of motion for this tachyonic model as:
$\frac{\ddot{\phi}}{(1-\dot{\phi}^{2})}+3H\dot{\phi}+\frac{1}{V}\frac{dV}{d\phi}=0.$ (12)
Now using Eq.(4) and Eq.(5); $\phi(t)$ and $V(\phi)$ can easily be written in the from of $H$ and $\dot{H}$ as following :
$\phi(t)=\int{\left(dt\sqrt{-\frac{2\dot{H}}{3H^{2}}}\right)}$ (13)
$V(t)=\frac{3H^{2}}{\kappa^{2}}\sqrt{1+\frac{2\dot{H}}{3H^{2}}}$ (14)
Now using above two equations we will find the potential and their respective inflection point behaviour for the three cases to be discussed below.
2.1 For Case-I: $a(t)=\alpha t^{n}$
In this subsection we consider the simplest model for describing the expansion of the universe in certain cosmological scenarios, namely, the power law form of scale factor suggesting that the universe’s expansion follows a simple power-law relationship with time. i.e. $a(t)=\alpha t^{n}$. Here $\alpha$ is a constant describing the overall scale of the universe and the exponent $n$ determines the rate of expansion and is greater than 00 for the universe to experience the accelerated expansion. For this kind of model the Hubble parameter is of the following form:
$H=\frac{n}{t}$ (15)
Now on substituting this value of $H$ in Eq.(14) and using Eq.(13) we found the following form of the potential:
$V(\phi)=V_{0}(\phi-\phi_{0})^{-2}$ (16)
here $V_{0}=\frac{2n}{\kappa^{2}}\sqrt{1-\frac{2}{3n}}$ and $n>2/3$ for the expansion of the universe.
Figure 1: The plot shows the variation of $\phi$ with potential $V(\phi)$ for different values of the exponent factor $n$ ($n=0.7$ (red solid line) and $n=0.9$ (blue dotted line))
From above Eq.(16) we found there is no such inflection point for this particular model and we can also see that from the Fig. (1) as well that the curve show no inflection but one can observe that the scalar field is rolling slowly along the potential. Now using Eq.(12) and on applying the slow roll approximation i.e. $\dot{\phi}\approx constant\implies\ddot{\phi}\approx 0$ we found the following evolution of scalar field with respect to time:
$\phi(t)=\sqrt{\frac{2}{3n}(t^{2}-t_{0}^{2})}+\phi_{0}$ (17)
here $t=t_{0}$ when $\phi=\phi_{0}$
Figure 2: The plot shows the evolution of $\phi$ with time $t$ for different values of the exponent factor $n$
Fig.(2) shows the evolution of $\phi$ for different values of $n.$
2.2 For Case-II: $a(t)=\gamma e^{\beta t}$
For this particular case we have taken a simplest exponential form of the scale factor $a(t)=\gamma e^{\beta t}$ where $\gamma$ and $\beta$ are the constants. $\beta$ determines the overall rate at which the scale factor changes and can have two possibilities for the values of first is $\beta>0$ implies the exponential growth of the universe i.e. expanding universe over time and the second possibility is $\beta<0$ implies the exponential decay i.e. contracting universe. As we know during early epoch the universe experiences expansion so we can discard the second possibility. Now the second constant $\gamma$ is a normalization factor that sets the overall scale of the universe at a particular time $t$ and $H$ for this exponentially expanding universe is proportional to the constant:
$H=\beta$ (18)
Now again using this value of $H$ and Eq.(13) we hot the following form of potential:
$V(\phi)=\frac{3\beta^{2}}{\kappa^{2}}$ (19)
therefore for this particular model there is no inflection point because the potential is coming out to be constant and is consistent with the case that of quintessence model. From Eq.(12), (19) and after applying the slow roll approximation the scalar field $\phi$ is coming out to be a constant and same result can be obtained from Eq.(10).
2.3 For Case-III: $a(t)=\eta t^{n}e^{\beta t}$
For this case we consider the universe in Quasi-exponential phase with the scale factor given by $a(t)=\eta t^{n}e^{\beta t}$ i.e. product of power law and exponential, here $\eta$, n and $\beta$ are the constants. For this quasi-exponential model the Hubble parameter is coming out to be:
$H=\frac{n}{t}+\beta$ (20)
Now using Eq.(13), (14) and (20), one can found the following form of potential in terms of $t$ given as follow:
$V(t)=\frac{3}{\kappa^{2}}\sqrt{\left(\frac{n+\beta t}{t}\right)^{4}-\frac{2n}{3}\left(\frac{n+\beta t}{t^{2}}\right)^{2}}$ (21)
and
$t=-\frac{n}{\beta}+\frac{n}{\beta}\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}$ (22)
On substituting the Eq. (22) in (21) we got the following form of potential $V(\phi)$:
$V(\phi)=\frac{2n}{\kappa^{2}(\phi-\phi_{0})^{2}}\sqrt{\left(1+\sqrt{\frac{3}{2n}}\beta(\phi-\phi_{0})\right)^{2}-\frac{2}{3n}}$ (23)
On solving the above equation the inflection points for this particular potential are: $\phi_{0}=\frac{3\beta\phi\pm 2+6\sqrt{6n}}{3\beta}$
On substituting the value of potential (23) in equation of motion for this quasi-exponential tachyonic model, we get the following form of the evolution of scalar field:
$\phi(t)=\sqrt{2-2\left(\frac{\exp{(\frac{n+\beta t}{3\beta^{2}}+\frac{2-(\phi-\phi_{0})^{2}}{2})}}{(n+\beta t)^{n/3\beta^{3}}}\right)^{2}}+\phi_{0}$ (24)
Figure 3: The plot shows the evolution of $\phi$ with time $t$ for the first inflection point with different values of the exponent factor $n$ and $\beta$
Figure 4: The plot shows the evolution of $\phi$ with time $t$ for the second inflection point for the different values of the exponent factor $n$ and $\beta$
3 Dynamical Analysis:
In this section we performs the dynamical analysis of our inflection point tachyon model. Dynamical analysis is the study of the evolution of the universe, its behaviour over the time and allow us to analyse the various components of the universe like dark matter, dark energy, ordinary matter. The background equations are rewritten as an independent system of first order ODEs so that the stability at equilibrium points may be examined. Setting the dimensionless variables as:
$X=\dot{\phi}$ (25a)
$Y=\frac{\kappa\sqrt{V}}{\sqrt{3}H}$ (25b)
$\Gamma=\frac{V(d^{2}V/d\phi^{2})}{(dV/d\phi)^{2}}$ (25c)
$\lambda=-\frac{dV/d\phi}{\kappa\sqrt{V^{3}}}$ (25d)
Here we have obtained an autonomous 2 dimensional system which has been studied by [59] and other two parameters ((25c)) and ((25d)) still has $\phi$ dependency and on differentiating above equations with respect to $\eta=loga$ we get the following set of 2 dimensional autonomous equations:
$X^{\prime}=\frac{dX}{d\eta}=(X^{2}-1)(3X-\lambda\sqrt{3}Y)$ (26a)
$Y^{\prime}=\frac{dY}{d\eta}=\frac{-Y}{2}\left(\frac{3Y^{2}}{\sqrt{1-X^{2}}}(w_{m}-X^{2}+1)-3(w_{m}+1)+\sqrt{3}\lambda XY\right)$ (26b)
$\lambda^{\prime}=\frac{d\lambda}{d\eta}=-\left(\frac{\sqrt{27}}{2}-\sqrt{3}\Gamma\right)XY\lambda^{2}$ (26c)
With a constraint equation:
$\frac{Y^{2}}{\sqrt{1-X^{2}}}+\Omega_{m}=1$ (27)
The equation of state and the energy density of the tachyonic scale field in terms of these dimensionless parameters can be written as:
$w_{\phi}=X^{2}-1$ (28)
$\Omega_{\phi}=\frac{Y^{2}}{\sqrt{1-X^{2}}}$ (29)
and the effective equation of motion can be written as:
$w_{eff}=w_{m}\left(1-\frac{Y^{2}}{\sqrt{1-X^{2}}}\right)-Y^{2}\sqrt{1-X^{2}}\$ (30)
The critical points for the Eq.((25a)) and ((25b)) are discussed below and are listed in the table(1):
$X$ $Y$ $w_{eff}$ $w_{\phi}$ $\Omega_{\phi}$ Stability
00 00 $w_{m}$ $-1$ 00 Saddle point
±$\pm 1$ 00 00 00 -- Unstable Nodes
$1/5$ $\frac{\sqrt{3}}{5\lambda}$- $\approx-0.96$ $\approx 1$ Stable node
Table 1: Above table shows the critical points of the system (25a)-(30) with stability properties.
For the inflection point there are four critical points as shown in fig(5) and are discussed below:
(a.) Point O: For this critical point $X=0$ and $Y=0$; the energy density of tachyonic scalar field is coming out to be 00 signifying the matter dominated epoch. From the Friedmann constraint ((27)) we can see that $\Omega_{m}=1$ and $w_{eff}=w_{m}$ again signifying the matter dominated epoch, thus this origin point of phase space refers to the saddle point having following eigenvalues:
$\nu_{1}=-3<0$ (31)
and
$\nu_{2}=\frac{3}{2}(w_{m}+1)>0$ (32)
This saddle point has an attractor behaviour towards X-axis as shown in fig(5) and at this point the value of $w_{\phi}=-1$ imitating the cosmological constant behaviour.
Figure 5: phase portait of the dynamical system with $\lambda=1$ and $w=0$ near inflection point. Here point $A$ in the phase space shows the region where universe undergoes accelerated expansion
Figure 6: shows the phase portait for the above mentioned set of dynamical equations with $\lambda=1$ and $w=0$.
(b.) Point B__\__ : Here for this point we got the following critical points $X=-1$ and $Y=0$. At this point the tachyonic equation of state vanishes i.e. $w_{\phi}=0$. At this critical point we have following eigenvalues:
$\nu_{1}=6>0$ (33a)
and
$\nu_{2}=\frac{3}{2}(w_{m}+1)>0$ (33b)
signifying this critical point B__\__ is an unstable node.
(c.) Point B+: (X=1,Y=0) this point also represent unstable node having same eigenvalues as mentioned above in Eq.33
(d.) Point A: $(X=\frac{1}{5},Y=\frac{\sqrt{3}}{5\lambda})$ is the stable node having eigen values:
$\nu_{1,2}=\frac{1}{2}\left(-4.8-1.9w_{m}\pm\sqrt{(0.7-4.3w_{m}-3.9w_{m}^{2})}\right)<0$ (34)
Fig.(3) and (4) shows the phase portait of the derived set of autonomous equations and the region near the point A depict the accelerated expansion of the universe.
4 Conclusion
In this paper the evolution of minimally coupled Tachyonic scalar field at the inflection point has been investigated in an accelerating universe. The inflection point signifies a transition between regions of stability and instability. Near this point, the field’s dynamics has been investigated for different expansion factors. We have mainly considered the three categories of scale (expansion) factor namely: (a) Power law expansion ($a(t)=\alpha t^{n}$), (b) Exponential form ($a(t)=\gamma e^{\beta t}$) and (c) Quasi-Exponential expansion ($a(t)=\eta t^{n}e^{\beta t}$); for all these categories, the potentials and the corresponding possible inflection points have been calculated and the dynamical system of equations for the model has been obtained. Fig.(3) and (4) represents the critical points and the stability of the derived dynamical system of equations and the region near point A shows the accelerated expansion of the universe.
In addition to what is being done, it is important to study the inflection point of scalar field models in the presence of quantum gravity effects. In particular, scalar field dynamics plays a significant role in the post bounce scenario of loop quantum cosmology as it gives rise to inflationary solution [64, 65]. We leave it as our future pursuit to investigate how inflection points are effected due to quantum geometric correction for scalar fields.
Acknowledgement
The research by M.K. was carried out in Southern Federal University with financial support from the Ministry of Science and Higher Education of the Russian Federation (State contract GZ0110/23-10-IF).